# Algorithm Design and Prototype Validation for Reconfigurable Intelligent Sensing Surface: Forward-Only Transmission

### v1

@@ I Introduction
Moreover, numerous studies have focused on sensing-assisted communication, exploring the potential for collaboration on channel estimation through sensing results or using sensing as an alternative to traditional channel estimation methods. This approach presents both new opportunities and challenges. Specifically, limited CSI is utilized for communication and rough sensing during the initial phase. The collected sensing information is then used to enhance both communication and sensing in the subsequent phase[22]. This two-phase ISAC transmission protocol for RIS-aided MIMO ISAC systems enables cooperative operation between the two functions. Similarly, [23] presents two deep learning networks designed for beam selection and power allocation, respectively. With the assistance of sensing data, these networks work together to maximize the communication channel capacity. Furthermore, [24] employs both active and passive elements for communication. Specifically, the active elements are utilized to sensing the directions of the BS and users in the uplink and downlink, respectively. The sensing results are then applied to facilitate uplink wireless information transfer (WIT) and downlink WET. Partial related works are presented in Table I.
@@
Sensing scheme Interference suppression Robust design Implementation
@@
RISS-based[24] DoA estimation
@@
  • We design a dual-functional RIS, which involving active elements and passive elements, termed the reconfigurable intelligent sensing surface (RISS). Thus, the RISS has the ability of sensing and signal processing from active elements. Leveraging this capability, the RISS eliminates the need for base station (BS) control and simplifies traditional CSI feedback by utilizing real-time sensing information. As a result, the proposed system can operate in a purely forward transmission process.
  • As the substitution of CSI, we use the sensing information including direction of arrival (DOA) estimation and identification recognition results for RISS passive beamforming design under near-field and far-field channel models. This approach not only enhances the target signal but also suppresses interference, leading to an improved signal-to-interference plus noise ratio (SINR) at the receiver.
  • To further evaluate performance, we developed the hardware for the RISS, which includes a 2-bit RIS with 100 passive elements and a uniformly linear array (ULA) with 4 active elements. We implemented two sensing algorithms—DOA estimation and RF identification recognition—as well as the communication system using software-defined radio (SDR). Extensive experiments were conducted to validate the performance of the proposed target enhancement and interference suppression schemes by leveraging the sensing results. The findings demonstrate its effectiveness and provide a strong foundation for future research on advanced communication schemes and sensing beamforming design.
@@ II System Model
Far-field and near-field channel models are both adopted for our system in this paper. We assume two $M$ antenna transmitters (Tx) transmit signal simultaneously to a single antenna receiver (Rx) via the RISS, which consists of total $(N_{a}+N)$ elements, including $N_{a}$ active elements and $N=N_{x}\times N_{y}$ passive elements, as depicts in Fig. 1. We assume these active elements involves the RISS have the ability of signal processing, e.g., DOA estimation and identification recognition. This enables the RISS to operate independently of the BS, eliminating the need for feedback control from the Rx (as shown in Fig. 1), and allows the passive phase shift matrix of RISS to be designed based on sensing information, without relying on traditional CSI feedback. Additionally, there is no direct link between the Rx and Tx.1 and Tx.2 due to blockages, without loss of generality.
@@ II-A Far-field Model
Quasi-static block line-of-sight (LoS) channels are taken into account. Specifically, the channel between the Tx-to-RISS and RISS-to-Rx are denoted as $\mathbf{G}_{i}\in\mathbb{C}^{N\times M},\forall i=\{1,2\}$ and $\mathbf{h}\in\mathbb{C}^{N\times 1}$, then we have
@@
${[\bm{\alpha}(\vartheta,\varphi)]_{n}=\frac{1}{\sqrt{N_{y}}}e^{\left(\bmod\left(n,N_{y}\right)-1\right)\mathrm{i}\varphi}\frac{1}{\sqrt{N_{x}}}e^{\left(\left\lfloor n/N_{x}+1\right\rfloor-1\right)\mathrm{i}\vartheta},}$ (3)
$[\beta(\varpi)]_{n}=\frac{1}{\sqrt{M}}e^{(n-1)\mathrm{i}\varpi},$ (4)
[…]
@@ II-B Near-field Model
Observe from Eq. (5) and Eq. (10) that the far-field and near-field models share similar expressions [^1] In practice, when the transmitter and receiver are sufficiently distant from the RISS, electromagnetic waves transition from near-field spherical wave modeling to far-field plane wave modeling, rendering the near-field and far-field models equivalent.. However, the specific channel expressions (i.e., Eq. (1)-(2) and Eq. (8)-(9)) differ. We will compare the disparities between the two channels through theoretical algorithm design and practical examination in the subsequent sections.
@@ III-A Far-field RISS-Aid Communications
where $P_{i}$ is the transmit power.
@@
Note that Eq. (11) can be divided into two sub-objective, result in
@@
To maximize the second objective, the maximal ratio transmit (MRT) technique is adopted. Thus, the proof is completed. ∎
@@
Inspired by [27], we utilize IRM algorithm for rank-one recovery. Specifically, we first obtain the solution of (P1.1) and denote it as $\mathcal{C}_{2}^{0}$. Once $\text{Rank}\left(\mathcal{C}_{2}^{0}\right)=1$, the optimal solution $\bm{\Theta}=\bm{\Theta}_{h}\bm{\Theta}_{G}^{0}$ is achieved for the original objective (P1). Otherwise, the objective (P1.1) transforms into
@@
$r\mathbf{I}_{N-1}-\mathbf{V}_{t}^{H}\bm{\mathcal{C}}_{2,t}\mathbf{V}_{t}\succeq\mathbf{0},$ (18c)
@@
where $\mathbf{V}_{t}$ denotes the matrix composed of eigenvectors corresponding to $[\lambda_{1},\ldots,\lambda_{N-1}]$, and $\lambda_{1}<\lambda_{2}<\ldots<\lambda_{N}$ represent the ascending eigenvalues of $\mathbf{C}_{2,t}$ in the $t^{\text{th}}$ iteration. Note that (P1.2) is a convex optimization problem, which can be solved by standard optimization method, such as the interior-point method. The IRM algorithm based solution is summarized in Algorithm 1.
@@
0:    Optimal solution $\mathcal{C}_{2}$ and optimal phase shift matrix $\bm{\Theta}$.
@@ III-B Near-field RISS-Aid Communications
$\varphi_{k}=2\pi\left(\frac{d_{\text{Tx2R}}^{2,k}+d_{\text{R2Rx}}^{k}}{\lambda}-\left\lfloor\frac{d_{\text{Tx2R}}^{2,k}+d_{\text{R2Rx}}^{k}}{\lambda}\right\rfloor\right)+\varpi,$ (25)
where $\varpi\in\left[0,2\pi\right]$ denotes the flexible variable shared among all passive elements of the RISS. The AO-based algorithm is summarized in Algorithm 2333Note that the input location information of Algorithm 2 can be directly estimated by the active elements of the RISS, which includes distance estimation [28, 29] and DOA estimation[30, 31, 32, 33]. In this paper, we assume that the distances between the Tx, Rx, and the center of the RISS are fixed, while varying the angles of incident signals..
@@ III-C Robust Design
We assume that the Tx, Rx and the RISS are all on the same plane, and the location of the Rx is perfectly known (i.e., we have the knowledge of $\mathbf{b}_{\text{Rx}}$), while the actual incident angle of the interference signal (i.e., $\phi^{\text{ele}}_{G,1}$) is within
$\phi^{\text{ele}}_{G,1}\in\left[\hat{\phi}{{}^{\text{ele}}_{G,1}}-\delta,\hat{\phi}{{}^{\text{ele}}_{G,1}}+\delta\right],$ (26)
where $\hat{\phi}{{}^{\text{ele}}_{G,1}}$ denotes the estimated incident angle of Tx. 1, and $\delta>0$ represents the boundary of the DOA estimation error. The fundamental robust design approach for both far-field and near-field scenarios involves dividing the interval $[\hat{\phi}{{}^{\text{ele}}_{G,1}}-\delta,\hat{\phi}{{}^{\text{ele}}_{G,1}}+\delta]$ into $L$ sub-intervals with a step size of $2\delta/{(L-1)}$, denoted by $\hat{\phi}{{}^{\S}_{1,l}},\forall l\in L$, and applying the suppression constraints to all angles within the interval.
@@ III-C1 Robust Design for Far-field Model
where $\bm{\mathcal{A}}_{1,l}=\bm{\alpha}(\hat{\vartheta}^{\S}_{1,l},\varphi_{G,1})\bm{\alpha}^{H}(\hat{\vartheta}^{\S}_{1,l},\varphi_{G,1})$, with $\hat{\vartheta}^{\S}_{1,l}=\pi\sin(\pi/2)\cos(\hat{\phi}{{}^{\S}_{1,l}})$ due to the assumption that all the Tx, Rx, and the RISS are on the same plane. The IRM algorithm can be continuously used to solve the objective (P1.3) for rank-one recovery.
@@ III-C2 Robust Design for Near-field Model
Similar to the far-field scenario, when the actual incident angle of the interference signal is within $\phi^{\text{ele}}_{G,1}\in[\hat{\phi}{{}^{\text{ele}}_{G,1}}-\delta,\hat{\phi}{{}^{\text{ele}}_{G,1}}+\delta]$, the location of Tx. 1 (i.e., $\mathbf{b}_{\text{Tx},1}$) and the distance between the $k^{\text{th}}$ RISS element and Tx. 1 vary, as depicted in Eq. (6). Thus, the objective (P2) can be rewritten as

### v2

@@ I Introduction
Moreover, numerous studies have focused on sensing-assisted communication, exploring the potential for collaboration on channel estimation through sensing results or using sensing as an alternative to traditional channel estimation methods. This approach presents both new opportunities and challenges. Specifically, limited CSI is utilized for communication and rough sensing during the initial phase. The collected sensing information is then used to enhance both communication and sensing in the subsequent phase[22]. This two-phase ISAC transmission protocol for RIS-aided MIMO ISAC systems enables cooperative operation between the two functions. Similarly, [23] presents two deep learning networks designed for beam selection and power allocation, respectively. With the assistance of sensing data, these networks work together to maximize the communication channel capacity.
Furthermore, a RIS structure combining active and passive elements, referred to as reconfigurable intelligent sensing surface (RISS), was proposed to enhance sensing capabilities in [24]. Building on this architecture, a joint time-of-arrival (TOA) and direction-of-arrival (DOA) estimation method for localization was developed, along with the derivation of the corresponding Cramér–Rao bound (CRB) [11], further demonstrating the potential of RISS in sensing tasks. In addition, [25] proposed a RISS-assisted wireless powered communication network (WPCN), where the active elements are employed to sensing the DOA of the BS and the users in the uplink and downlink, respectively. The sensing results are then utilized to facilitate uplink wireless information transfer (WIT) and downlink WET. Similarly, [26] introduced a multi-RISS-assisted ISAC framework, which jointly optimizes transmit and reflective beamforming to minimize the worst-case CRB for target estimation while guaranteeing communication quality. More recently, [27] developed a space-time-coding metasurface that is capable of real-time incident DOA sensing. The proposed coding matrix dynamically adapts the beam direction, enabling coordinated and real-time operation of communication and sensing functions. Partial related works are presented in Table I.
@@
Sensing scheme Comm. Interf. suppr. Robust design Implementation
@@
RISS-sensing[24] DOA estimation
RISS-SWIPT[25] DOA estimation
RISS-location[11] DOA estimation
RISS-ISAC[26] DOA estimation
STCM[27] DOA estimation
@@
  • We extend the conventional sensing paradigm of RISS by proposing a dual-sensing enhanced communication architecture. In this framework, active sensing elements within the RISS are utilized for both DOA estimation and RF identification, enabling effective discrimination between target and interference signals.
  • Based on the acquired sensing information, we design the reflection phase shift matrix for the passive elements. Both near-field and far-field communication channel models are considered. Furthermore, robust beamforming strategies are developed to enhance the target signal and suppress interference, taking into account potential sensing errors, thereby significantly improving the signal-to-interference-plus-noise ratio (SINR) at the receiver.
  • By leveraging the real-time sensing capability of RISS, the proposed system eliminates the reliance on base station (BS) control and reduces the complexity of traditional CSI feedback. Consequently, the RISS can be directly integrated into existing communication systems and operate in a purely forward transmission manner, which simplifies the overall communication process.
  • To validate the proposed framework, a prototype hardware system is developed, consisting of a 2-bit RIS with 100 passive elements and a uniform linear array (ULA) with 4 active elements. We implement two key sensing algorithms—DOA estimation and RF identification—along with the communication module on a software-defined radio (SDR) platform. Extensive experiments are conducted to evaluate the performance of the proposed target enhancement and interference suppression schemes. The results confirm the effectiveness of our approach and lay a solid foundation for future research on ISAC beamforming and system design.
@@ II System Model
Far-field and near-field channel models are both adopted for our system in this paper. We assume two $M$ antennas transmitters (Tx) transmit signal simultaneously to a single antenna receiver (Rx) via the RISS, which consists of total $(N_{a}+N)$ elements, including $N_{a}$ active elements and $N=N_{x}\times N_{y}$ passive elements, as depicts in Fig. 1. We assume these active elements involves the RISS have the ability of signal processing, e.g., DOA estimation and identification recognition. This enables the RISS to operate independently of the BS, eliminating the need for feedback control from the Rx (as shown in Fig. 1), and allows the passive phase shift matrix of RISS to be designed based on sensing information, without relying on traditional CSI feedback. Additionally, there is no direct link between the Rx and Tx.1 and Tx.2 due to blockages, without loss of generality [^1] In fact, the orthogonal polarization between the incident and reflected signals of the RIS effectively blocks the direct link between the Tx and Rx in Section IV..
Remark 1
The placement of both active and passive elements in the RISS plays a crucial role in the proposed framework. In this work, the active elements are arranged in a ULA (or an L-array for two-dimensional DOA estimation), along one side of the RISS, as illustrated in Fig. 1. This configuration is adopted for the following reasons: 1. DOA estimation algorithms based on ULA and L-arrays have been extensively studied and validated. 2. The passive elements retain a complete uniform planar array (UPA) structure, which allows for the direct application of conventional RIS beamforming and optimization methods. 3. The spatial separation between active and passive elements helps mitigate potential hardware coupling effects.
Moreover, we advocate for a moderate increase in the number of active elements in the RISS, rather than an unbounded expansion. This recommendation is supported by two key observations. First, the reflected beam inherently has a finite beamwidth, which provides robustness to DOA estimation errors and reduces the system’s sensitivity to sensing inaccuracies. Second, a larger number of passive elements provides greater degrees of freedom for interference suppression, thereby enhancing system robustness against angular deviations in sensing.
Additionally, increasing the number of active elements incurs significant hardware costs, including more RF chains and higher power consumption, making large-scale deployment impractical in many scenarios. Therefore, a balanced design should aim to ensure sufficient sensing capability with a limited number of active elements, while maximizing communication performance through an expanded passive array.
The relationship between sensing accuracy and communication performance has been analyzed in [25], offering further insights into the influence of sensing precision on system performance.
@@ II-A Far-field Model
Quasi-static block line-of-sight (LoS) channels are taken into account [^2] It is worth noting that the proposed scheme for target signal enhancement and interference suppression based on the LoS channel can be directly extended to more complex Rician channels. As shown in [25], in scenarios where the LoS component is predominant, relying solely on LoS-based sensing enables near-optimal communication performance relative to full CSI, with substantially lower pilot overhead.. Specifically, the channel between the Tx-to-RISS and RISS-to-Rx are denoted as $\mathbf{G}_{i}\in\mathbb{C}^{N\times M},\forall i=\{1,2\}$ and $\mathbf{h}\in\mathbb{C}^{N\times 1}$, then we have
@@
$\bm{\beta}(\varpi_{G,i})=\frac{1}{\sqrt{M}}\left[1,e^{\mathrm{i}\varpi_{G,i}},e^{\mathrm{i}2\varpi_{G,i}},\cdots,e^{\mathrm{i}(M-1)\varpi_{G,i}}\right],$ (3)
$\bm{\alpha}(\vartheta_{X},\varphi_{X})=\bm{\alpha}_{x}\left(\vartheta_{X}\right)\otimes\bm{\alpha}_{y}\left(\varphi_{X}\right),X\in\left\{h,(G,i)\right\}$ (4)
where $\vartheta_{X}=2\pi d/\lambda\cos(\phi_{X}^{\text{ele}})=\pi\cos(\phi_{X}^{\text{ele}})$ by setting $d/\lambda=1/2$ without loss of generality [^3] Excessively large $d$ can introduce grating lobes, leading to energy dispersion and distortion of the radiation pattern, while excessively small $d$ may cause mutual coupling between antennas. Considering these factors, setting $d=\lambda/2$ is a widely accepted compromise that balances both effects.. $d$ and $\lambda$ represent the element spacing and carrier wavelength, respectively. Similarly, we express $\varphi_{X}=2\pi d/\lambda\sin(\phi^{\text{ele}}_{X})\cos(\phi^{\text{azi}}_{X})=\pi\sin(\phi^{\text{ele}}_{X})\cos(\phi^{\text{azi}}_{X})$ and $\varpi_{G,i}=\pi\sin(\phi_{G,i}^{\text{dep}})$[28, 24]. And we have
$\bm{\alpha}_{x}(\vartheta_{X})=\frac{1}{\sqrt{N_{x}}}\left[1,e^{\mathrm{i}\vartheta_{X}},\cdots,e^{\mathrm{i}(N_{x}-1)\vartheta_{X}}\right],$
$\bm{\alpha}_{y}(\varphi_{X})=\frac{1}{\sqrt{N_{y}}}\left[1,e^{\mathrm{i}\varphi_{X}},\cdots,e^{\mathrm{i}(N_{y}-1)\varphi_{X}}\right],$ (5)
Additionally, $\phi^{\text{azi}}_{G,i}$ ($\phi^{\text{ele}}_{G,i}$) and $\phi^{\text{dep}}_{G,i}$ denote the azimuth (elevation) angle of arrival (AOA) and the angle of departure (AOD) from the transmitter to the RISS (Tx-to-RISS). Furthermore, $\phi^{\text{azi}}_{h}$ ($\phi^{\text{ele}}_{h}$) represent the azimuth (elevation) angle of departure (AOD) from the RISS to the receiver (RISS-to-Rx), respectively, as depicts in the left-bottom of Fig. 1. Thus, the received signal at the Rx can be formulated as
@@ II-B Near-field Model
Observe from Eq. (6) and Eq. (11) that the far-field and near-field models share similar expressions [^4] In practice, when the Tx and Rx are sufficiently far from the RIS, the waves transition from a near-field spherical wavefronts to far-field planar wavefronts, making the near-field and far-field models effectively equivalent. Generally, the far-field assumption holds when $R>\frac{2D^{2}}{\lambda}$. Here, $D$ represents the maximum dimension of the antenna, typically its diameter or maximum length. $R$ and $\lambda$ denote the distance between Tx/Rx and RIS and wavelength, respectively.. However, the specific channel expressions (i.e., Eq. (1)-(2) and Eq. (9)-(10)) differ. We will compare the disparities between the two channels through theoretical algorithm design and practical examination in the subsequent sections.
@@ III-A Far-field RISS-Aid Communications
where $P_{i}$ is the transmit power and |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm.
@@
Note that Eq. (12) can be decomposed into two sub-objectives, resulting in
@@
It is important to emphasize that our objective is to adjust the reflection matrix $\bm{\Theta}$ of the RISS through a sensing-assisted scheme, allowing the RISS to independently control the signals in free space. And from the Tx’s perspective, the focus is solely on maximizing the term $\bm{\beta}^{T}(\varpi_{G,i})\mathbf{v}_{i}$.
According to the Cauchy-Schwarz inequality, the maximum of $|\bm{\beta}^{T}(\varpi_{G,i})\mathbf{v}_{i}|$ is achieved when $\mathbf{v}_{i}$ is aligned with the conjugate of $\bm{\beta}(\varpi_{G,i})$, i.e., $\mathbf{v}_{i}=\bm{\beta}^{\dagger}(\varpi_{G,i})$. This can then be normalized as Eq. (13), which is consistent with the maximal ratio transmission (MRT) technique. ∎
@@
Inspired by [31], we utilize IRM algorithm for rank-one recovery. Specifically, we first obtain the solution of (P1.1) and denote it as $\bm{\mathcal{C}}_{2}^{0}$. Once $\text{Rank}\left(\bm{\mathcal{C}}_{2}^{0}\right)=1$, the optimal solution $\bm{\Theta}=\bm{\Theta}_{h}\bm{\Theta}_{G}^{0}$ is achieved for the original objective (P1). Otherwise, the objective (P1.1) transforms into
@@
$r\mathbf{I}_{N-1}-\mathbf{V}_{t}^{H}\bm{\mathcal{C}}_{2}^{t}\mathbf{V}_{t}\succeq\mathbf{0},$ (19c)
@@
where $\mathbf{V}_{t}$ denotes the matrix composed of eigenvectors corresponding to $[\lambda_{1},\ldots,\lambda_{N-1}]$, and $\lambda_{1}<\lambda_{2}<\ldots<\lambda_{N}$ represent the ascending eigenvalues of $\bm{\mathcal{C}}_{2}^{t}$ in the $t^{\text{th}}$ iteration. $r$ is the positive relaxation variable. As $r$ approaches zero, the undesired eigenvalue $[\lambda_{1},\cdots,\lambda_{N-1}]$ are forced to become zeros, which implies that $\bm{\mathcal{C}}_{2}^{t}$ satisfies the rank-one constraint. $\epsilon_{l}$ and $\varsigma$ represent gradually increasing auxiliary variables that control the rate at which $r$ decreases. Note that (P1.2) is a convex optimization problem, which can be solved by standard optimization method, such as the interior-point method. The IRM algorithm based solution is summarized in Algorithm 1.
@@
0:    Optimal solution $\bm{\mathcal{C}}_{2}^{t}$ and optimal phase shift matrix $\bm{\Theta}$.
@@ III-B Near-field RISS-Aid Communications
$\varphi_{k}=2\pi\left(\frac{d_{\text{Tx2R}}^{2,k}+d_{\text{R2Rx}}^{k}}{\lambda}-\left\lfloor\frac{d_{\text{Tx2R}}^{2,k}+d_{\text{R2Rx}}^{k}}{\lambda}\right\rfloor\right)+\psi,$ (26)
where $\psi\in\left[0,2\pi\right]$ denotes the flexible variable shared among all passive elements of the RISS. The AO-based algorithm is summarized in Algorithm 2666Note that the input location information of Algorithm 2 can be directly estimated by the active elements of the RISS, which includes distance estimation [32, 33] and DOA estimation[34, 35, 36, 37]. In this paper, we assume that the distances between the Tx, Rx, and the center of the RISS are fixed, while varying the angles of incident signals..
@@ III-C Robust Design
We assume that the Tx, Rx and the RISS are all on the same plane, and the location of the Rx is perfectly known (i.e., we have the knowledge of $\mathbf{b}_{\text{Rx}}$), while the actual incident angle of the interference signal (i.e., $\phi^{\text{azi}}_{G,1}$) is within
$\phi^{\text{azi}}_{G,1}\in\left[\hat{\phi}{{}^{\text{azi}}_{G,1}}-\delta,\hat{\phi}{{}^{\text{azi}}_{G,1}}+\delta\right],$ (27)
where $\hat{\phi}{{}^{\text{azi}}_{G,1}}$ denotes the estimated incident angle of Tx. 1, and $\delta>0$ represents the boundary of the DOA estimation error. The fundamental robust design approach for both far-field and near-field scenarios involves dividing the interval $[\hat{\phi}{{}^{\text{azi}}_{G,1}}-\delta,\hat{\phi}{{}^{\text{azi}}_{G,1}}+\delta]$ into $L$ sub-intervals with a step size of $2\delta/{(L-1)}$, denoted by $\hat{\phi}{{}^{\S}_{1,l}},\forall l\in L$, and applying the suppression constraints to all angles within the interval.
@@ III-C1 Robust Design for Far-field Model
where $\bm{\mathcal{A}}_{1,l}=\bm{\alpha}(\hat{\vartheta}^{\S}_{1,l},\varphi_{G,1})\bm{\alpha}^{H}(\hat{\vartheta}^{\S}_{1,l},\varphi_{G,1})$, with $\hat{\vartheta}^{\S}_{1,l}=\pi\cos(\pi/2)=0$ due to the assumption that all the Tx, Rx, and the RISS are on the same plane. The IRM algorithm can be continuously used to solve the objective (P1.3) for rank-one recovery.
@@ III-C2 Robust Design for Near-field Model
Similar to the far-field scenario, when the actual incident angle of the interference signal is within $\phi^{\text{azi}}_{G,1}\in[\hat{\phi}{{}^{\text{azi}}_{G,1}}-\delta,\hat{\phi}{{}^{\text{azi}}_{G,1}}+\delta]$, the location of Tx. 1 (i.e., $\mathbf{b}_{\text{Tx},1}$) and the distance between the $k^{\text{th}}$ RISS element and Tx. 1 vary, as depicted in Eq. (7). Thus, the objective (P2) can be rewritten as
@@
III-D Computational Complexity and Convergence Analysis
This section analyzes the computational complexity of Algorithm 1 and Algorithm 1. Specifically, the complexity of solving each optimization problem (P1.2) using the interior-point method is $\mathcal{O}\left(N^{3.5}\log\frac{1}{\zeta}\right)$, where $\zeta$ denotes the prescribed accuracy [38]. Let $I_{\mathrm{IRM}}$ denote the number of iterations required by the IRM algorithm, where each iteration involves an eigenvalue decomposition, incurring a computational cost of $\mathcal{O}(N^{3})$. Consequently, the total computational complexity of Algorithm 1 can be expressed as $\mathcal{O}\left(I_{\mathrm{IRM}}\left(N^{3.5}\log\frac{1}{\zeta}+N^{3}\right)\right).$
In contrast, solving the optimization problem (P2) requires an exhaustive search over $\Delta_{\varphi}N$ phase shift combinations, where each search involves recalculating $S_{2}-\eta S_{1}$, introducing a complexity of $\mathcal{O}(N+\Delta_{\phi}N)$. Here, $\Delta_{\varphi}$ and $\Delta_{\phi}$ represent the phase shift resolution and the robust design resolution, respectively. Thus, the total computational complexity of Algorithm 2 with $I_{\mathrm{AO}}$ iterations is approximately $\mathcal{O}\left(I_{\mathrm{AO}}\Delta_{\varphi}N\left(N+\Delta_{\phi}N\right)\right)\approx\mathcal{O}\left(I_{\mathrm{AO}}\Delta_{\varphi}\Delta_{\phi}N^{2}\right).$
(a) Convergence of Algorithm 1.
(b) Convergence of Algorithm 2.
Figure 3: Convergence of both algorithm, where $N=100$, $\Delta_{\varphi}=4$ for 2-bit RISS, and $\Delta_{\phi}=2\delta/0.1$.
The convergence behavior of both algorithms is illustrated in Fig. 3. It is observed that both algorithms converge within a finite number of iterations. Furthermore, the number of required iterations increases when the interference suppression constraints become more stringent, e.g., $\tau=-20$ dB and $\eta=5$. Therefore, appropriate interference suppression strategies can be selected based on the practical scenario to balance the trade-off between computational efficiency and system performance.